# Jet launching from the Kerr black hole magnetosphere: An electrogeodesic approach

### v1

@@ Abstract
The launch of relativistic jets of plasma on astrophysical to cosmological scales are observed in a variety of astrophysical sources, from active galactic nuclei to X-rays binaries. While these jets can be reproduced by the general relativistic magneto-hydrodynamic (GRMHD) and particle-in-cells (GRPIC) simulations of the dynamical Kerr magnetosphere, the development of analytic models to describe the physics of the jets has remained limited. A key challenge is to analytically describe the individual trajectories of accelerated charged particles which ultimately build up the jet and emit radiation. In this work, we provide a first simple but fully analytical model of jet launching from the Kerr magnetosphere based on the motion of charged particles. To that end, we use the integrability of electrogeodesic motion in the Kerr monopole magnetosphere to study the ejection of charged particles near the poles. This enables us to derive (i) a criterion for the rotation axis to constitute a stable latitunal equilibrium position, thereby representing an idealized jet, (ii) the expression for the magnetic frame-dragging effect, and (iii) the condition for an asymptotic observer to measure blueshifted particles emanating from the black hole surroundings. Our study reveals that particles can be accelerated only in a specific region whose maximal radius depends on the spin and magnetization of the black hole. Alongside these results, we provide a detailed review of the construction of test magnetospheres from (explicit and hidden) symmetries of the Kerr geometry and the condition for the separability of the electrogeodesic motion in a test magnetosphere, which serves as a basis for the model we study in this work.
@@ Introduction
Relativistic jets of plasma are observed in a variety of astrophysical systems, from active galactic nuclei (AGN) containing a supermassive black hole (SMBH) in its centter to stellar mass black hole involved in X-rays binaries [1, 2]. They can extend up to thousand of light years, therefore playing a major role in galaxy evolution by redistributing matter on large scales [3]. These jets are believed to be powered by highly magnetized spinning black holes which accrete plasma and accelerate particles at very high energy. Yet the precise mechanisms at play to produce and sustain these powerful jets, together with a detailed understanding of their structure, still remains to be understood [4]. While catalogue of astrophysical systems powering jets have been monitored during the last decades [5, 6], the new born horizon-scale electromagnetic astronomy initiated by the Event Horizon Telescope [7, 8] now enables us to image relativistic jets of various systems with an unprecedented resolution, among which the M87 SMBH [9]. These observation coupled to a better characterization of the properties of the jet could provide a new window to further constrain the mass and spin of SMBH [10].
@@
To a large extent, our theoretical understanding of energy extraction from highly magnetized spinning black hole relies on the seminal Blandford-Znajek (BZ) model [11]. The BZ solution describes a slowly rotating monopole magnetosphere surrounding the Kerr black hole, which is sourced by a dilute plasma. This Kerr magnetosphere successfully generates outgoing Poynting fluxes triggered by the toroidal components of the magnetic field. The BZ model relies on the force-free (FF) regime where the electromagnetic field satisfies $E\cdot B=0$ and $E^{2}<B^{2}$ everywhere. The FF approximation ensures that the plasma density $\rho$ is sufficient to source the magnetosphere but small enough to keep it in a high magnetization regime, i.e. $\sigma\sim B^{2}/\rho\gg 1$. See [12, 13, 14, 15, 16, 17] for more details on the BZ process. While the validity of the FF approximation and the asymptotic stability of the BZ solution casted some doubts in the early days [18], the progress in general relativistic (GR) magneto-hydrodynamics (MHD) simulations of dynamical black hole magnetospheres [19, 20, 21] allowed to address these questions and eventually numerically reproduce the final steady-state BZ solution and its Poynting winds [22, 23, 24, 25]. See [26] for a pedagogical review on the theoretical description of force-free magnetosphere.
@@
Nevertheless, the BZ model and its associated GRMHD simulations suffer from several limitations to properly characterize the jet launching mechanism. First, the BZ solution is perturbative in the spin and thus cannot account for high spinning black hole. So far, no non-perturbative extension has been found [27, 28, 29, 30]. Second, and more importantly, the GRMHD simulations rely on a fluid description of the plasma which is inherently incomplete. In particular, it fails to track the motion of individual charged particles in the magnetosphere which requires a kinetic approach. Switching from the fluid-based to a particle-based description of the jet is crucial to make progress. Moreover, the FF condition forbids the formation of pairs and electric gaps which are believed to play an important role in the jet dynamics. Even more importantly, the FF condition prevents charged particle to locally accelerate, which happens only when $E\cdot B\neq 0$, i.e. in the presence of non-ideal electric fields. The developments of GR/Particle-In-Cells (PIC) simulations have played a major role in addressing these difficulties [31, 32, 33]. These simulations, based on the ZELTRON code, allow us to follow the individual particles trajectories of the plasma evolving in a dynamical magnetosphere where the force free condition can be violated. While they could also reproduce jet launching from Kerr immersed in a asymptotic uniform field [34], they also revealed new effects such as the build up of current sheet at the equator where magnetic reconnection appears to play a crucial role to eject energy from the black hole. See [35, 36, 37, 38, 39, 40, 41] for further results along this line. Besides their qualitative results, these simulations underline the crucial need to properly characterize the motion of individual charged particles (as opposed to fluid) in the black hole magnetosphere to properly account of the complicated mechanisms at play in the accretion/ejection paradigm.
@@
The important progresses in the development of the GR/PIC simulations of the black hole magnetosphere dramatically contrast with the analytic investigations, which have remained scarce [42, 43]. Indeed, the inherent complexity in solving the electrogeodesic motion around magnetized black hole has limited most studies to a mixture of analytical and numerical treatments [44, 45, 46, 47, 48, 49, 50, 51, 52, 53]. This prevents us to obtain close analytic predictions for the structure of the jet, its opening angle, the size of the acceleration and escape zones or the azimuthal velocity of charged particles in the different regions of the jet, which should play a key role to describe the emitted radiation. In contrast, the neutral geodesic motion in the Kerr geometry is well known to be fully integrable analytically [54]. The existence of a hidden symmetry, leading to the famous Carter constant, allows one to fully separate the radial and polar motion, which in turn, allows one to separate the time and azimuthal motion. In general, this two step process breaks down when a test-field magnetosphere is turned on, such that the electrogeodesic equations are no longer separable. This is typically the case of the well-known and extensively studied Wald solution [66], which describes an asymptotic uniform magnetic field around a Kerr black hole. However, the situation is not completely hopeless.
@@
Indeed, as initially shown in the seminal work of Carter [54], the separability of the geodesic motion on the Kerr black hole extends to the electrogeodesic motion on the electrically and magnetically charged Kerr-Newman (KN) black hole, i.e. the electrovacuum solution to the Einstein-Maxwell equations extending the Kerr solution and parametrized by four constants: the mass $M$, the spin $a$, the electric charge $\mathcal{Q}$} and the magnetic charge $\mathcal{P}$}. To our knowledge, the only systematic effort to extend Carter’s analytic result and classify the electrogeodesic motion in the KN geometry was presented much later by Hackman and Xu in [55]. See [56, 57, 59, 58] for related but more specific investigations. At first, it would appear natural to use these elegant results to study the jet formation in the KN magnetosphere. However, the presence of an electric charge $\mathcal{Q}$} makes the model not so realistic, since the surrounding plasma would rapidly discharge the black hole. As for the magnetic charge, this limitation does not exist given that no elementary particles carrying magnetic charge are known to exist. Yet, the magnetic charge $\mathcal{P}$} of the black hole is expected to be negligible compared to its mass, i.e. ${\mathcal{P}}/M\ll 1$, so that gravity remains the dominant effect. This more realistic regime, i.e. ${\mathcal{Q}}=0$ and ${\mathcal{P}}/M\ll 1$, corresponds to the test field approximation of the magnetic KN geometry. Interestingly, taking this limit reduces the problem to the magnetically charged, vacuum magnetosphere on the Kerr black hole which, although well known, has not received much attention in the literature. Yet, it appears to be the simplest analytic starting point to model the a rotating black hole vacuum magnetosphere. At this level, a remarkable observation is that the separability of the electrogeodesic motion in the KN black hole remains valid in the test field approximation. This allows one to obtain closed form solutions to the equations of motion of charged particles. In this work, we use this path to analytically study the acceleration and the ejection of test charged particles around a magnetized Kerr black hole. Surprisingly, even in this simple vacuum magnetosphere, the exact electrogeodesic solutions reveal a rich behavior and allows one to identify a subset of trajectories that build up a jet of accelerated particles. Let us summarize the new qualitative results for the physics of the jet emerging from this study:
@@
  • We present the analytic expression for the $4$-velocity $u^{\mu}\partial_{\mu}$ [Eq. (4.6) below] and 4-acceleration $a^{\mu}\partial_{\mu}$ [Eq. (4.7)] of charged particles in the whole black hole exterior, which allows us to contemplate several key properties of the system. Contrary to the azimuthal and polar acceleration $(a^{\theta},a^{\varphi}$), the radial acceleration $a^{r}$ is shown to be proportional to both the magnetic charge $\mathcal{P}$} (carried by the magnetosphere) and the black hole spin $a$, such that it vanishes in the non-rotating and/or non-magnetized case. This demonstrates the key role played by the presence of a spinning magnetically charged magnetosphere in the ejection mechanism. Second, the polar acceleration turns out to be maximal at the equator and vanishing at the pole. In contrast, the radial acceleration vanishes at the equator and becomes maximal at the poles, suggesting the build up of jet of charged particles preferentially around the poles. For negative charged particles, $\kappa<0$, the acceleration is outwards and the particles are ejected if ${\mathcal{P}}>0$.
@@
  • To understand the structure of the ejection at the poles, and thus of the jet, we further analyze the stable and unstable equilibrium positions of the polar motion depending on the energy and angular momentum $(\epsilon,\ell)$ of the charged particle. The presence of the magnetosphere dramatically changes the properties of the polar potential. The difference with the geodesic motion is best parametrized by a new parameter $\beta=-\kappa{\mathcal{P}}/a$ where $\kappa$ is the charge-to-mass ratio of the particle. First, particles can reach the Northern rotation axis $\theta=0$ only for $\ell=a\beta$. Focusing on this family of electrogeodesics, we identify a new condition [Eq. (4.22) below] for the Northern rotation axis to be a stable latitudinal position, namely the energy of the particle must be either smaller or greater than a threshold depending on $\beta$. Instead, trajectories violating this condition admit a stable latitudinal position at some $\theta_{\ast}\in\;]0,\pi/2[$.
@@
  • We also study the fate of the frame-dragging effect in the Kerr monopole. Using the expression of the azimuthal angular velocity [Eq. (4.27) below], we show that the magnetic monopole introduces a new magnetic frame-dragging which completely dominates the gravitational frame-dragging. Indeed, we find that the former decays as $1/r^{2}$ while the later decays as $1/r^{3}$, showing that the new magnetic frame-dragging impacts charged particles on larger scales. Furthermore, in contrast to the gravitational frame-dragging which forces the particles to rotate in a prograde trajectory, the magnetic counterpart forces particles to have prograde or retrograde motion depending on their charge, i.e. on $\kappa$.
@@
Since this investigation is made possible thanks to the separability of the electrogeodesic motion in this specific model of the Kerr magnetosphere, it is interesting to understand the fundamental reasons why this elegant integrability occurs. While this is well known for the Kerr-Newman geometry, it is interesting to revisit this question from the point of view of the test field approximation. Indeed, other exact solutions such as the Wald solution which has been extensively studied fails to have separable electrogeodesic motion. To make this work as self-contained as possible, we present, prior to our study of charged particle ejection, a review of (i) the construction of exact solutions of the Maxwell equations in the Kerr geometry using the Killing symmetries and (ii) a detailed derivation of the geometrical conditions for preserving the separability, i.e. the Carter constant, for the electrogeodesic motion in a given test field magnetosphere. We also include a detailed derivation of the separable electrogeodesic equations in the Kerr magnetic monopole magnetosphere. This review allows us to present, besides the well known Wald construction, other results which have not received the same focus, such as the construction of exact vacuum solutions of Maxwell equations using Killing-Yano tensors known as the Penrose currents. We also use this opportunity to introduce new results, such as a new construction to build an exact non-vacuum solution of the Maxwell equation using Killing vectors, leading to a new exact Maxwell solution on the Kerr geometry [Eqs. (2.42)–(2.43) below]. In the end, this review provides the necessary technical tools to appreciate the fundamental reasons ensuring the separability of the electrogeodesic motion in the Kerr monopole magnetosphere, which descends from the properties of the Killing-Yano tensor of the Kerr geometry.
This work is organized as follows. In Section 2, we present the detailed review of the construction of exact test magnetospheres models from the explicit and hidden Killing symmetries, i.e. using the Killing vectors and the Killing-Yano tensors and apply them to the Kerr geometry. We present an alternative to the Wald construction in Section 2.1.3. Section 3 is devoted to the review on the separability of the electrogeodesic motion. We first review how the standard conditions for the geodesics are modified for the electrogeodesic case. Then, we derive the geometric condition for preserving the Carter constant. The detailed demonstration of the separability in the Kerr monopole magnetosphere is rederived in Section 3.4, together with the integral solutions for the electrogeodesic. In Section 4, we apply this result to study the ejection of charged particles from the Kerr magnetosphere and the structure of the jet. Finally, Section 5 provides a discussion of our results and the open directions to explore.
@@
In this section, we review the key constructions allowing one to obtain exact vacuum solutions of Maxwell equations using the symmetries of the spacetime. Consider therefore the Maxwell equations in a curved spacetime which take the form
$\nabla_{\mu}F^{\mu\nu}$ $=J^{\mu}$ (2.1)
$\nabla_{\mu}F_{\nu\alpha}+\nabla_{\nu}F_{\alpha\mu}+\nabla_{\alpha}F_{\mu\nu}$ $=0$ (2.2)
$\nabla_{\mu}J^{\mu}$ $=0$ (2.3)
where the Faraday tensor $F_{\mu\nu}$ can be decomposed in term of the gauge-potential $A=A_{\mu}\textrm{d}x^{\mu}$ as
$F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ (2.4)
@@
In the following, we shall review how to construct specific solutions of these equations. The focus will be on the Kerr metric which reads
@@
Consider a Killing vector $k^{\mu}\partial_{\mu}$ defined by $\nabla_{(\mu}k_{\nu)}=0$. This definition implies that it also satisfies the following equations
@@
Writing down the source-free ($J^{\mu}=0$) Maxwell equation (2.1) in terms of the potential $A_{\mu}\textrm{d}x^{\mu}$, one finds
@@
$\xi^{\mu}\partial_{\mu}=\partial_{t}\qquad\text{and}\qquad\chi^{\mu}\partial_{\mu}=\partial_{\varphi},$ (2.11)
which naturally gives rises to a two parameters family of source-free Maxwell solutions of the form
$A^{(\alpha,\beta)}_{\mu}\textrm{d}x^{\mu}=\alpha\xi_{\mu}\textrm{d}x^{\mu}+\beta\chi_{\mu}\textrm{d}x^{\mu}.$ (2.12)
@@
such that $\mathcal{Q}$} encodes the strength of the field (actually the total electric charge, as we shall see below). Equation (2.12) then yields
$A^{(\alpha,0)}_{\mu}\textrm{d}x^{\mu}=\alpha\xi_{\mu}\textrm{d}x^{\mu}=\frac{{\mathcal{Q}}}{2M}\left(1-\frac{2Mr}{\Sigma}\right)\textrm{d}t+\frac{a{\mathcal{Q}}r\sin^{2}{\theta}}{\Sigma}\textrm{d}\varphi.$ (2.17)
@@
${A^{\prime}}^{(\alpha,0)}_{\mu}\textrm{d}x^{\mu}=-\frac{{\mathcal{Q}}r}{\Sigma}\textrm{d}t+\frac{a{\mathcal{Q}}r\sin^{2}{\theta}}{\Sigma}\textrm{d}\varphi.$ (2.18)
The associated Faraday tensor $F^{(\alpha,0)}=\textrm{d}A^{(\alpha,0)}=\textrm{d}{A^{\prime}}^{(\alpha,0)}$ is given terms of the Carter coframe (2.15) by
$\boxed{F^{(\alpha,0)}=2\alpha\nabla_{\mu}\xi_{\nu}\;\textrm{d}x^{\mu}\wedge\textrm{d}x^{\nu}=-\frac{{\mathcal{Q}}(r^{2}-a^{2}\cos^{2}{\theta})}{\Sigma^{2}}\,e^{0}\wedge e^{1}+\frac{2a{\mathcal{Q}}r\cos{\theta}}{\Sigma^{2}}\,e^{2}\wedge e^{3}}$ (2.19)
@@
which corresponds to an electric monopole. While this is the simplest solution one can construct, its electric charge makes it a less appealing model for the magnetosphere. Indeed, the plasma around an electrically charged magnetosphere would quickly discharge any charge excess so that the electrically neutral configuration is the natural assumption when considering the black hole magnetosphere on long timescale. In order to consider a magnetosphere model without electric charge, one can either (i) combine it to the solution $F^{(0,\beta)}$ generated by the azimuthal Killing vector and appropriately choose the combination to make the total electric charge vanishes, this is the well known Wald construction [66], or (ii) one can consider its dual solution $\ast F^{(\alpha,0)}$, which instead has zero electric charge but a non-vanishing magnetic charge.
@@
$\tilde{A}^{(\alpha,0)}=\tilde{A}^{(\alpha,0)}_{\mu}\textrm{d}x^{\mu}=\frac{{\mathcal{P}}\cos{\theta}}{\Sigma}\left(a\textrm{d}t-(r^{2}+a^{2})\textrm{d}\varphi\right).$ (2.25)
@@
Let us consider the second branch of solutions which allows one to construct asymptotically uniform magnetic field following Wald’s construction [66]. The second branch corresponds to $\alpha=0$ in Eq. (2.12) and we introduce the constant parameter $\mathcal{B}$} such that

### v2

# arXiv:2601.05048v2[astro-ph.HE] 02 Mar 2026

@@ Abstract
The launch of relativistic jets of plasma on astrophysical to cosmological scales is observed in a variety of astrophysical sources, from active galactic nuclei to X-ray binaries. While these jets can be reproduced by general relativistic magneto-hydrodynamics (GRMHD) and particle-in-cells (GRPIC) simulations of the dynamical Kerr magnetosphere, the development of analytic models to describe the physics of the jets has remained limited. A key challenge is to analytically describe the individual trajectories of accelerated charged particles, which ultimately build up the jet and emit radiation. In this work, we provide a first simple but fully analytical model of jet launching from the Kerr magnetosphere based on the motion of charged particles. To that end, we use the integrability of electrogeodesic motion in the Kerr monopole magnetosphere to study the ejection of charged particles near the poles. This enables us to derive (i) a criterion for the rotation axis to constitute a stable latitudinal equilibrium position, thereby representing an idealized jet, (ii) the expression for the magnetic frame-dragging effect, and (iii) the condition for an asymptotic observer to measure blueshifted particles emanating from the black hole surroundings. Our study reveals that particles can be accelerated only in a specific region whose maximal radius depends on the spin and magnetization of the black hole. Alongside these results, we provide a detailed review of the construction of test magnetospheres from (explicit and hidden) symmetries of the Kerr geometry and the condition for the separability of the electrogeodesic motion in a test magnetosphere, which serves as a basis for the model studied in this work.
@@ Introduction
Relativistic jets of plasma are observed in a variety of astrophysical systems, from active galactic nuclei (AGN) containing a supermassive black hole (SMBH) in its centter to stellar mass black hole involved in X-ray binaries [1, 2]. They can extend up to thousands of light years, therefore playing a major role in galaxy evolution by redistributing matter on large scales [3]. These jets are believed to be powered by highly magnetized spinning black holes which accrete plasma and accelerate particles at very high energy. Yet the precise mechanisms at play to produce and sustain these powerful jets, together with a detailed understanding of their structure, still remains to be understood [4]. While catalogs of astrophysical systems powering jets have been monitored during the last decades [5, 6], the newborn horizon-scale electromagnetic astronomy initiated by the Event Horizon Telescope [7, 8] now enables us to image relativistic jets of various systems with an unprecedented resolution, among which the M87 SMBH [9]. These observations coupled to a better characterization of the properties of the jet could provide a new window to further constrain the mass and spin of SMBHs [10].
@@
To a large extent, our theoretical understanding of energy extraction from highly magnetized spinning black hole relies on the seminal Blandford-Znajek (BZ) model [11]. The BZ solution describes a slowly rotating monopole magnetosphere surrounding the Kerr black hole, which is sourced by a dilute plasma. This Kerr magnetosphere successfully generates outgoing Poynting fluxes triggered by the toroidal components of the magnetic field. The BZ model relies on the force-free (FF) regime where the electromagnetic field satisfies $E\cdot B=0$ and $E^{2}<B^{2}$ everywhere. The FF approximation ensures that the plasma density $\rho$ is sufficient to source the magnetosphere but small enough to keep it in a high magnetization regime, i.e. $\sigma\sim B^{2}/\rho\gg 1$. See [12, 13, 14, 15, 16, 17] for more details on the BZ process. While the validity of the FF approximation and the asymptotic stability of the BZ solution casted some doubts in the early days [18], the progress in general relativistic (GR) magneto-hydrodynamics (MHD) simulations of dynamical black hole magnetospheres [19, 20, 21] allowed to address these questions and eventually numerically reproduce the final steady-state BZ solution and its Poynting winds [22, 23, 24, 25]. See [26] for a pedagogical review on the theoretical description of force-free magnetospheres.
@@
Nevertheless, the BZ model and its associated GRMHD simulations suffer from several limitations to properly characterize the jet launching mechanism. First, the BZ solution is perturbative in the black hole spin and thus cannot account for rapidly spinning black holes. So far, no non-perturbative extension has been found [27, 28, 29, 30]. Second, and more importantly, the GRMHD simulations rely on a fluid description of the plasma which is inherently incomplete. In particular, it fails to track the motion of individual charged particles in the magnetosphere, which requires a kinetic approach. Switching from the fluid-based to a particle-based description of the jet is crucial to make progress. Moreover, the FF condition forbids the formation of pairs and electric gaps which are believed to play an important role in the jet dynamics. Even more importantly, the FF condition prevents charged particles to locally accelerate, which happens only when $E\cdot B\neq 0$, i.e. in the presence of non-ideal electric fields. The developments of GR/Particle-In-Cell (PIC) simulations have played a major role in addressing these difficulties [31, 32, 33]. These simulations, based on the ZELTRON code, allow us to follow the individual particle trajectories of the plasma evolving in a dynamical magnetosphere where the force free condition can be violated. While they could also reproduce jet launching from a Kerr black hole immersed in an asymptotic uniform field [34], they also revealed new effects such as the build up of current sheet at the equator where magnetic reconnection appears to play a crucial role to eject energy from the black hole. See [35, 36, 37, 38, 39, 40, 41] for further results along this line. Besides their qualitative results, these simulations underline the crucial need to properly characterize the motion of individual charged particles (as opposed to fluid) in the black hole magnetosphere to properly account of the complicated mechanisms at play in the accretion/ejection paradigm.
@@
The important progresses in the development of the GR/PIC simulations of the black hole magnetosphere dramatically contrast with the analytic investigations, which have remained scarce [42, 43]. Indeed, the inherent complexity in solving the electrogeodesic motion around magnetized black holes has limited most studies to a mixture of analytical and numerical treatments [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. This prevents us to obtain close analytic predictions for the structure of the jet, its opening angle, the size of the acceleration and escape zones or the azimuthal velocity of charged particles in the different regions of the jet, which should play a key role to describe the emitted radiation. In contrast, the neutral geodesic motion in the Kerr geometry is well known to be fully integrable analytically [55]. The existence of a hidden symmetry, leading to the famous Carter constant, allows one to fully separate the radial and polar motion, which in turn, allows one to separate the time and azimuthal motion. In general, this two step process breaks down when a test-field magnetosphere is turned on, such that the electrogeodesic equations are no longer separable. This is typically the case of the well-known and extensively studied Wald solution [68], which describes an asymptotic uniform magnetic field around a Kerr black hole. However, the situation is not completely hopeless.
@@
Indeed, as initially shown in the seminal work of Carter [55], the separability of the geodesic motion on the Kerr black hole extends to the electrogeodesic motion on the electrically and magnetically charged Kerr-Newman (KN) black hole, i.e. the electrovacuum solution to the Einstein-Maxwell equations extending the Kerr solution and parametrized by four constants: the mass $M$, the spin parameter $a$, the electric charge $\mathcal{Q}$} and the magnetic charge $\mathcal{P}$}. To our knowledge, the only systematic effort to extend Carter’s analytic result and classify the electrogeodesic motion in the KN geometry was presented much later by Hackman and Xu in [56]. See [57, 58, 60, 59] for related but more specific investigations. At first, it would appear natural to use these elegant results to study the jet formation in the KN magnetosphere. However, the presence of an electric charge $\mathcal{Q}$} makes the model not so realistic, since the surrounding plasma would rapidly discharge the black hole. As for the magnetic charge, this limitation does not exist given that no elementary particles carrying magnetic charge are known to exist. Yet, the magnetic charge $\mathcal{P}$} of the black hole is expected to be negligible compared to its mass, i.e. ${\mathcal{P}}/M\ll 1$, so that gravity remains the dominant effect. This more realistic regime, i.e. ${\mathcal{Q}}=0$ and ${\mathcal{P}}/M\ll 1$, corresponds to the test field approximation of the magnetic KN geometry. Interestingly, taking this limit reduces the problem to the magnetically charged, vacuum magnetosphere on the Kerr black hole which, although well known, has not received much attention in the literature. Yet, it appears to be the simplest analytic starting point to model the a rotating black hole vacuum magnetosphere. At this level, a remarkable observation is that the separability of the electrogeodesic motion in the KN black hole remains valid in the test field approximation. This allows one to obtain closed form solutions to the equations of motion of charged particles. In this work, we use this path to analytically study the acceleration and the ejection of test charged particles around a magnetized Kerr black hole. Surprisingly, even in this simple vacuum magnetosphere, the exact electrogeodesic solutions reveal a rich behavior and allows one to identify a subset of trajectories that build up a jet of accelerated particles. Let us summarize the new qualitative results for the physics of the jet emerging from this study:
@@
  • We present the analytic expression for the $4$-velocity $u^{\mu}$ [Eq. (4.6) below] and 4-acceleration $a^{\mu}$ [Eq. (4.7)] of charged particles in the whole black hole exterior, which allows us to contemplate several key properties of the system. Contrary to the azimuthal and polar acceleration $(a^{\theta},a^{\varphi}$), the radial acceleration $a^{r}$ is shown to be proportional to both the magnetic charge $\mathcal{P}$} (carried by the magnetosphere) and the black hole spin $a$, such that it vanishes in the non-rotating and/or non-magnetized case. This demonstrates the key role played by the presence of a spinning magnetically charged magnetosphere in the ejection mechanism. Second, the polar acceleration turns out to be maximal at the equator and vanishing at the pole. In contrast, the radial acceleration vanishes at the equator and becomes maximal at the poles, suggesting the build up of jet of charged particles preferentially around the poles. For negative charged particles, the acceleration is outwards and the particles are ejected if ${\mathcal{P}}>0$.
@@
  • To understand the structure of the ejection at the poles, and thus of the jet, we further analyze the stable and unstable equilibrium positions of the polar motion depending on the specific energy and angular momentum $(\varepsilon,\ell)$ of the charged particle. The presence of the magnetosphere dramatically changes the properties of the polar potential. The difference with the geodesic motion is best parametrized by a new parameter $\beta=-\kappa{\mathcal{P}}/a$ where $\kappa$ is the charge-to-mass ratio of the particle. First, particles can reach the Northern rotation axis $\theta=0$ only for $\ell=a\beta$. Focusing on this family of electrogeodesics, we identify a new condition [Eq. (4.22) below] for the Northern rotation axis to be a stable latitudinal position, namely the energy of the particle must be either smaller or greater than a threshold depending on $\beta$. Instead, trajectories violating this condition admit a stable latitudinal position at some $\theta_{\ast}\in\;]0,\pi/2[$.
@@
  • We also study the fate of the frame-dragging effect in the Kerr monopole. Using the expression of the azimuthal angular velocity [Eq. (4.27) below], we show that the magnetic monopole introduces a new magnetic frame-dragging which completely dominates the gravitational frame-dragging. Indeed, we find that the former decays as $1/r^{2}$ while the later decays as $1/r^{3}$, showing that the new magnetic frame-dragging impacts charged particles on larger scales. Furthermore, in contrast to the gravitational frame-dragging which forces particles to rotate in a prograde trajectory, the magnetic counterpart forces particles to have prograde or retrograde motion depending on their charge, i.e. on $\kappa$. This magnetic frame dragging is expected to have interesting consequences on the precession of a misaligned jet of charged particles [61].
@@
Since this investigation is made possible thanks to the separability of the electrogeodesic motion in this specific model of the Kerr magnetosphere, it is interesting to understand the fundamental reasons why this elegant integrability occurs. While this is well known for the Kerr-Newman geometry, it is interesting to revisit this question from the point of view of the test field approximation. Indeed, other exact solutions such as the Wald solution, which has been extensively studied, fail to have separable electrogeodesic motion. To make this work as self-contained as possible, we present, prior to our study of charged particle ejection, a review of (i) the construction of exact solutions of the Maxwell equations in the Kerr geometry using the Killing symmetries and (ii) a detailed derivation of the geometrical conditions for preserving the separability, i.e. the Carter constant, for the electrogeodesic motion in a given test field magnetosphere. We also include a detailed derivation of the separable electrogeodesic equations in the Kerr magnetic monopole magnetosphere. This review allows us to present, besides the well known Wald construction, other results which have not received the same focus, such as the construction of exact vacuum solutions of Maxwell equations using Killing-Yano tensors known as the Penrose currents. We also use this opportunity to introduce new results, such as a non-vacuum solution of the Maxwell equation in the Kerr geometry built upon the two main Killing vectors [Eqs. (2.41)–(2.42) below]. In the end, this review provides the necessary technical tools to appreciate the fundamental reasons ensuring the separability of the electrogeodesic motion in the Kerr monopole magnetosphere, which descends from the properties of the Killing-Yano tensor of the Kerr geometry.
This article is organized as follows. In Section 2, we present a detailed review of the construction of exact test magnetosphere models from the explicit and hidden Killing symmetries, i.e. using the Killing vectors and the Killing-Yano tensors and apply them to the Kerr geometry. We present an alternative to the Wald construction in Section 2.1.3. Section 3 is devoted to the review on the separability of the electrogeodesic motion. We first review how the standard conditions for the geodesics are modified for the electrogeodesic case. Then, we derive the geometric condition for preserving the Carter constant. The detailed demonstration of the separability in the Kerr monopole magnetosphere is rederived in Section 3.4, together with the integral solutions for the electrogeodesic. In Section 4, we apply this result to study the ejection of charged particles from the Kerr magnetosphere and the structure of the jet. Finally, Section 5 provides a discussion of our results, as well as of the open directions to explore.
@@
In this section, we review the key constructions allowing one to obtain exact vacuum solutions of Maxwell equations using the symmetries of the spacetime. Consider therefore the Maxwell equations in a curved spacetime; they take the form
$\nabla_{\nu}F^{\mu\nu}$ $=4\pi J^{\mu}$ (2.1)
$\nabla_{\mu}F_{\nu\alpha}+\nabla_{\nu}F_{\alpha\mu}+\nabla_{\alpha}F_{\mu\nu}$ $=0,$ (2.2)
where the Faraday tensor $F_{\mu\nu}$ can be expressed in term of the gauge-potential $A=A_{\mu}\textrm{d}x^{\mu}$ as
$F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}.$ (2.3)
@@
In the following, we shall review how to construct specific solutions of these equations. The focus will be on the Kerr metric, which reads
@@
Consider a Killing vector $k^{\mu}$ on spacetime. By definition, it obeys $\nabla_{(\mu}k_{\nu)}=0$, which implies that it also satisfies the following equations:
@@
Writing down the source-free ($J^{\mu}=0$) Maxwell equation (2.1) in terms of the potential $A_{\mu}$, one finds
@@
$\xi=\partial_{t}\qquad\text{and}\qquad\chi=\partial_{\varphi},$ (2.10)
which naturally give rise to a two parameters family of source-free Maxwell solutions of the form
$A^{(\alpha,\beta)}_{\mu}=\alpha\xi_{\mu}+\beta\chi_{\mu}.$ (2.11)
@@
such that $\mathcal{Q}$} encodes the strength of the field (actually the total electric charge, as we shall see below). Equation (2.11) reduces then to $A^{(\alpha,0)}_{\mu}=\alpha\xi_{\mu}$, which yields
$A^{(\alpha,0)}=\frac{{\mathcal{Q}}}{2M}\left(1-\frac{2Mr}{\Sigma}\right)\textrm{d}t+\frac{a{\mathcal{Q}}r\sin^{2}{\theta}}{\Sigma}\textrm{d}\varphi.$ (2.16)
@@
${A^{\prime}}^{(\alpha,0)}=-\frac{{\mathcal{Q}}r}{\Sigma}\textrm{d}t+\frac{a{\mathcal{Q}}r\sin^{2}{\theta}}{\Sigma}\textrm{d}\varphi.$ (2.17)
The associated Faraday tensor $F^{(\alpha,0)}=\textrm{d}A^{(\alpha,0)}=\textrm{d}{A^{\prime}}^{(\alpha,0)}$ is given in terms of the Carter coframe (2.14) by
$\boxed{F^{(\alpha,0)}=-\frac{{\mathcal{Q}}(r^{2}-a^{2}\cos^{2}{\theta})}{\Sigma^{2}}\,e^{0}\wedge e^{1}+\frac{2a{\mathcal{Q}}r\cos{\theta}}{\Sigma^{2}}\,e^{2}\wedge e^{3}}$ (2.18)
@@
which corresponds to an electric monopole. While this is the simplest solution one can construct, its electric charge makes it a less appealing model for the magnetosphere. Indeed, the plasma around an electrically charged magnetosphere would quickly discharge any charge excess so that the electrically neutral configuration is the natural assumption when considering the black hole magnetosphere on a long timescale. In order to consider a magnetosphere model without electric charge, one can either (i) combine it to the solution $F^{(0,\beta)}$ generated by the azimuthal Killing vector and appropriately choose the combination to make the total electric charge vanishes, this is the well known Wald construction [68], or (ii) one can consider its dual solution $\ast F^{(\alpha,0)}$, which instead has zero electric charge but a non-vanishing magnetic charge.
@@
$\tilde{A}^{(\alpha,0)}=\frac{{\mathcal{P}}\cos{\theta}}{\Sigma}\left(a\textrm{d}t-(r^{2}+a^{2})\textrm{d}\varphi\right).$ (2.24)
@@
Let us consider the second branch of solutions (2.11), which leads to uniform magnetic fields following Wald’s construction [68]. The second branch corresponds to $\alpha=0$ in Eq. (2.11) and we introduce the constant parameter $\mathcal{B}$} such that